# Constructions of binary self-orthogonal singly-even minimal linear codes violating the Aschikhmin-Barg condition with few weights

### v2

@@ 2-A Basic knowledge of (vectorial) Boolean functions
A well-known class of bent functions is the Maiorana-McFarland (MM for short) class $\mathcal{M}$ [dillon1974elementary, mcfarland1973family], which gives by far the largest class of bent functions defined by the set of all the Boolean functions on ${\mathbb{F}}_{2}^{m}=\left\{(y,x):y,x\in{\mathbb{F}}_{2}^{t}\right\}$ with $m=2t$, of the form:
$f(y,x)=y\cdot\pi(x)+g(x),$ (2)
where $\pi$ is any permutation over ${\mathbb{F}}_{2}^{t}$ and $g$ is any Boolean function on ${\mathbb{F}}_{2}^{t}$. The dual function of the MM class $f$ is
$f^{*}(y,x)=x\cdot\pi^{-1}(y)+g(\pi^{-1}(y)).$ (3)
@@ 4 Constructions of binary SO singly-even minimal linear codes from Boolean functions
Since for all $b\in{\mathbb{F}}_{2}^{k}$, $W_{g}(x)\equiv 0\pmod{4}$, we can see that for all $(b,c_{1},c_{2})\in{\mathbb{F}}_{2}^{m}$,
@@ 5 Constructions of binary SO singly-even minimal linear codes from vectorial Boolean functions
    for any $u\in\mathbb{F}_{2^{k-1}}^{*}$ and any $b_{1},b_{2}\in\mathbb{F}_{2^{m}}$, it holds that $(-1)^{i}W_{f}(b_{1})+(-1)^{j}W_{G}(u,b_{2})+(-1)^{l}W_{H_{u}}(b_{1}+b_{2})\neq 2^{n}$, where $i,j,l\in\{0,1\}$ such that exactly two of them are 0;
@@
    If $f^{*}_{1}(0)=f^{*}_{2}(0)=f^{*}_{3}(0)=1$ and $f^{*}_{1}+f^{*}_{2}\neq f^{*}_{3}$, then the function $f=f_{1}f_{2}$ satisfies the conditions $(\mathbf{c}_{1})$ and $(\mathbf{c}_{2})$ in Definition 5.1.
@@
For any given $b\in\mathbb{F}_{2^{m}}$, we have
$W_{f}(b)$ $=$ $\sum\limits_{x\in\mathbb{F}_{2^{m}}}(-1)^{f_{1}(x)f_{2}(x)+\mathrm{Tr}_{2^{m}}(bx)}$ (16)
We first prove that the statement (1) holds. It follows from the assumption that Eq. (16) can be written as
@@
[…]
We now prove that the statement (2) holds. Combining Eq. (16) and the proof of the statement (1), we can obtain that if $f^{*}_{1}+f^{*}_{2}=f^{*}_{3}$, then
$W_{f}(b)$ $=$ $\left\{\begin{array}[]{ll}2^{m-1}+2^{t-1},&\mathrm{if}~b=0,\\ -3\cdot 2^{t-1},&\mathrm{if}~b\neq 0,(f^{*}_{1}(b),f^{*}_{2}(b))=(1,1),\\ 2^{t-1},&\mathrm{if}~b\neq 0,(f^{*}_{1}(b),f^{*}_{2}(b))\in\{(\mathrm{i}+1,\mathrm{i}),(0,0)\},\end{array}\right$. (26)
where $\mathrm{i}\in\{0,1\}$ and if $f^{*}_{1}+f^{*}_{2}\neq f^{*}_{3}$, then
$W_{f}(b)$ $=$ $\left\{\begin{array}[]{ll}2^{m-1}+2^{t-1},&\mathrm{if}~b=0,\\ (-1)^{\mathrm{i}}2^{t-1},&\mathrm{if}~b\neq 0,(f^{*}_{1}(b),f^{*}_{2}(b),f^{*}_{3}(b))\in\{(\mathrm{i},\mathrm{i},\mathrm{i}),\\ &(1+\mathrm{i},\mathrm{i},1+\mathrm{i}),(\mathrm{i},1+\mathrm{i},1+\mathrm{i})\}\\ (-1)^{\mathrm{i}}3\cdot 2^{t-1},&\mathrm{if}~b\neq 0,(f^{*}_{1}(b),f^{*}_{2}(b),f^{*}_{3}(b))=(\mathrm{i},\mathrm{i},1+\mathrm{i}),\end{array}\right$. (31)
where $\mathrm{i}\in\{0,1\}$.
Finally, incorporated with Eqs. (26) and (31), it is straightforward to obtain that $f$ satisfies the conditions $(\mathbf{c}_{1})$ and $(\mathbf{c}_{2})$ in Definition 5.1. ∎
@@
We now provide explicit constructions based on Proposition 5.3 and the well-known MM class.
Let $\lambda_{1},\lambda_{2}$ be distinct elements in $\mathbb{F}_{2^{t}}^{*}$, and $w_{1},w_{2}\in\mathbb{F}^{*}_{2^{t}}$ with $w_{1}\neq w_{2}$. For any $(y,x)\in\mathbb{F}_{2^{t}}\times\mathbb{F}_{2^{t}}$, we define the new Boolean functions as follows
$f(y,x)=$ $\mathrm{Tr}_{2^{t}}(\lambda_{1}y\pi(x))\mathrm{Tr}_{2^{t}}(\lambda_{2}y\pi(x)),$ (32)
$G(y,x)=$ $\mathrm{Tr}_{2^{t}}(\lambda_{3}y\pi(x))+((x+w_{1})^{2^{t}-1}+(x+w_{2})^{2^{t}-1})(y^{2^{t}-1}+1),$ (33)
$F(y,x)=$ $(f(y,x),G(y,x)),$ (34)
where $\lambda_{3}=\lambda_{1}+\lambda_{2}$, $\pi$ is a permutation on ${\mathbb{F}}_{2^{t}}$.
To simplify subsequent calculations, we introduce the following definition.
@@
$\mathcal{A}_{i}=\mathrm{Tr}_{2^{t}}(\beta_{2}\pi^{-1}(\lambda_{i}^{-1}\beta_{1}))$, where $\lambda_{3}=\lambda_{1}+\lambda_{2}$ and $1\leq i\leq 3$. Define $\mathcal{V}=\mathrm{Span}(\delta_{1},\delta_{2},\ldots,\delta_{t-2})$, where $\delta_{1},\delta_{2},\ldots,\delta_{t-2}$ are linearly independent over $\mathbb{F}_{2^{m}}$, and for any $x\in\mathcal{V}$, the equation $\mathrm{Tr}_{2^{t}}(w_{1}x)=\mathrm{Tr}_{2^{t}}(w_{2}x)=0$ holds.
@@
[berlekamp1967solution] Let $t$ be a positive integer and $a,b\in\mathbb{F}_{2^{t}}$ with $a\neq 0$. Then the quadratic equation $x^{2}+ax+b=0$ has solutions in $\mathbb{F}_{2^{t}}$ if and only if $\mathrm{Tr}_{2^{t}}(\frac{b}{a^{2}})=0$.
Before presenting the next result, we first note that, from Lemma 5.6, if $(\lambda_{1}+\lambda_{2})^{-1}=\lambda_{1}^{-1}+\lambda_{2}^{-1}$, then $t\geq 2$; otherwise, $t\geq 3$. Moreover, we adopt the notation $(\mathcal{V}\times\mathbb{F}_{2^{t}})^{*}=\mathcal{V}\times\mathbb{F}_{2^{t}}\setminus\{(0,0)\}$.
In the following, we first prove that the above $f$ and $G$ satisfy the conditions in Definition 5.1.
@@
Let $f$ and $G$ be defined as in Eqs. (32) and ($3$). Then $f$ and $G$ satisfy the conditions in Definition 5.1, where $m=2t,k=2$.
@@
On one hand, note that $f=f_{1}f_{2}$, where $f_{1}=\mathrm{Tr}_{2^{t}}(\lambda_{1}y\pi(x))$ and $f_{2}=\mathrm{Tr}_{2^{t}}(\lambda_{2}y\pi(x))$ belongs to the MM class. By Eq. (3), we have that $f_{1}^{*}={\rm Tr}_{2^{t}}(\lambda_{1}^{-1}x\pi^{-1}(y))$ and $f_{2}^{*}={\rm Tr}_{2^{t}}(\lambda_{2}^{-1}x\pi^{-1}(y))$. It follows from Proposition 5.3 that the Boolean function $f$ defined in Eq. (32) satisfies the conditions $(\mathbf{c}_{1})$ and $(\mathbf{c}_{2})$ in Definition 5.1.
@@
[…]
Thus $G(0,0)=0$ and it is clear that $vf+uG$ is not affine for any nonzero $(v,u)\in{\mathbb{F}}_{2}\times{\mathbb{F}}_{2}$. From Eq. (43) and Lemma 2.8, it follows that $\mathcal{C}_{G}$ is a minimal code. ∎
We now consider the explicit constructions of binary linear codes. After plugging the vectorial Boolean function $F$ defined as in (34) into the linear code ${\mathcal{C}}_{F}$ derived in (4), we can find that the resulting linear codes are SO and doubly-even. To obtain singly-even linear codes, we make a slight modification as follows.
${\mathcal{C}}_{F}=\left\{{\bf u}_{a,b}=((v,u)\cdot F(y,x)+b_{1}y+b_{2}x)_{(y,x)\in{\mathbb{F}}_{2}^{t}\times{\mathbb{F}}_{2}^{t}\backslash\{(\mathbf{0},\mathbf{0})\}}:(v,u)\in{\mathbb{F}}_{2}\times{\mathbb{F}}_{2},(b_{1},b_{2})\in\mathcal{V}\times{\mathbb{F}}_{2^{t}}\right\}.$ (35)
That is to say, our code ${\mathcal{C}}_{F}$ is a subspace of that derived in (4). Obviously, when the $b$ in Lemmas 3.6 and 5.2 belongs to a subspace of $\mathbb{F}_{2^{m}}$, the conclusions of these two lemmas still hold.
To compute the weight distribution of ${\mathcal{C}}_{F}$ derived in (35), it suffices to compute the values of the Walsh transform of $W_{F}$.

### v3

@@ 2-A Basic knowledge of (vectorial) Boolean functions
Remark 2.2.
For any bent function $f$ from ${\mathbb{F}}_{2}^{m}$ to ${\mathbb{F}}_{2}$ , let $\bar{f}=f+1$. Then it is trivial that for any $b\in{\mathbb{F}}_{2^{m}}$, $W_{\bar{f}}(b)=-W_{f}(b)$, implying that $\bar{f}$ is also bent and $\bar{f}^{*}=-f^{*}$.
A well-known class of bent functions is the Maiorana-McFarland (MM for short) class $\mathcal{M}$ [dillon1974elementary, mcfarland1973family], which gives by far the largest class of bent functions defined by the set of all the Boolean functions on ${\mathbb{F}}_{2}^{m}=\left\{(x,y):x,y\in{\mathbb{F}}_{2}^{t}\right\}$ with $m=2t$, of the form:
$f(x,y)=x\cdot\pi(y)+g(y),$ (2)
where $\pi$ is any permutation on ${\mathbb{F}}_{2}^{t}$ and $g$ is any Boolean function on ${\mathbb{F}}_{2}^{t}$. The dual function of the MM class $f$ is
$f^{*}(x,y)=y\cdot\pi^{-1}(x)+g(\pi^{-1}(x)).$ (3)
In the final of this subsection, we provide some formulas about computing the Walsh transform of some Boolean functions from Bent functions, which will be used in Section 5.
Lemma 2.3.
Let $f_{1}$ and $f_{2}$ be two different Bent functions from ${\mathbb{F}}_{2}^{m}$ to ${\mathbb{F}}_{2}$ with $m=2t$. Let $f=f_{1}f_{2}$. Then
$W_{f}(b)=\left\{\begin{array}[]{ll}2^{m-1}+2^{t-1}\left((-1)^{f^{*}_{1}(0)}+(-1)^{f^{*}_{2}(0)}-(-1)^{f^{*}_{3}(0)}\right),&\mathrm{if}~b=0,\\ 2^{t-1}\left((-1)^{f^{*}_{1}(b)}+(-1)^{f^{*}_{2}(b)}-(-1)^{f^{*}_{3}(b)}\right),&\mathrm{if}~b\neq 0,\end{array}\right$.
where $f_{3}=f_{1}+f_{2}$.
Proof.
For any given $b\in{\mathbb{F}}_{2^{m}}$, we have
$W_{f}(b)$ $=$ $\sum\limits_{x\in\mathbb{F}_{2^{m}}}(-1)^{f_{1}(x)f_{2}(x)+\mathrm{Tr}_{2^{m}}(bx)}$
By the above lemma, we can easily get the following corollary.
Corollary 2.4.
Let $f_{1}$ and $f_{2}$ be two different Bent functions from ${\mathbb{F}}_{2}^{m}$ to ${\mathbb{F}}_{2}$ with $m=2t$. Let $f=f_{1}f_{2}+f_{1}+f_{2}$. Then
$W_{f}(b)=\left\{\begin{array}[]{ll}-2^{m-1}+2^{t-1}\left((-1)^{f^{*}_{1}(0)}+(-1)^{f^{*}_{2}(0)}+(-1)^{f^{*}_{3}(0)}\right),&\mathrm{if}~b=0,\\ 2^{t-1}\left((-1)^{f^{*}_{1}(b)}+(-1)^{f^{*}_{2}(b)}+(-1)^{f^{*}_{3}(b)}\right),&\mathrm{if}~b\neq 0,\end{array}\right$.
where $f_{3}=f_{1}+f_{2}$.
Proof.
In this corollary, $f=(f_{1}+1)(f_{2}+1)+1$. Then the formula of the Walsh transform of $f$ can be directly obtained by Lemma 2.3 and Remark 2.2. ∎
@@ 4 Constructions of binary SO singly-even minimal linear codes from Boolean functions
Since for all $b\in{\mathbb{F}}_{2}^{k}$, $W_{g}(b)\equiv 0\pmod{4}$, we can see that for all $(b,c_{1},c_{2})\in{\mathbb{F}}_{2}^{m}$,
@@ 5 Constructions of binary SO singly-even minimal linear codes from vectorial Boolean functions
    for any $u\in\mathbb{F}_{2^{k-1}}^{*}$ and any $b_{1},b_{2}\in\mathbb{F}_{2^{m}}$, it holds that $(-1)^{i}W_{f}(b_{1})+(-1)^{j}W_{G}(u,b_{2})+(-1)^{l}W_{H_{u}}(b_{1}+b_{2})\neq 2^{m}$, where $i,j,l\in\{0,1\}$ such that exactly two of them are 0;
@@
    If $f^{*}_{1}(0)=f^{*}_{2}(0)=f^{*}_{3}(0)=1$ and there exist some elements $b\in{\mathbb{F}}_{2^{m}}^{*}$ such that $f^{*}_{1}(b)=1,f^{*}_{2}(b)=1,f^{*}_{3}(b)=0$, then the function $f=f_{1}f_{2}$ satisfies the conditions $(\mathbf{c}_{1})$ and $(\mathbf{c}_{2})$ in Definition 5.1.
@@
By Lemma 2.3, for any $b\in{\mathbb{F}}_{2^{m}}$, we have
$W_{f}(b)$ $=$ $\left\{\begin{array}[]{ll}2^{m-1}+2^{t-1}\left((-1)^{f^{*}_{1}(0)}+(-1)^{f^{*}_{2}(0)}-(-1)^{f^{*}_{3}(0)}\right),&\mathrm{if}~b=0,\\ 2^{t-1}\left((-1)^{f^{*}_{1}(b)}+(-1)^{f^{*}_{2}(b)}-(-1)^{f^{*}_{3}(b)}\right),&\mathrm{if}~b\neq 0.\end{array}\right$. (17)
We first prove the statement (1). It follows from the assumption that Eq. (17) can be written as
@@
By Eq. (23), it is easy to see that $f=f_{1}f_{2}$ satisfies the conditions $(\mathbf{c}_{1})$ and $(\mathbf{c}_{2})$ in Definition 5.1 under the assumption that $f^{*}_{1}(0)=f^{*}_{2}(0)=f^{*}_{3}(0)=1$ and and there exist some elements $b\in{\mathbb{F}}_{2^{m}}^{*}$ such that $f^{*}_{1}(b)=1,f^{*}_{2}(b)=1,f^{*}_{3}(b)=0$.
We are now proceed to prove the statement (2). Combining Eq. (17), we have
$W_{f}(b)$ $=$ $\left\{\begin{array}[]{ll}2^{m-1}+2^{t-1},&\mathrm{if}~b=0,\\ (-1)^{\mathrm{i}}2^{t-1},&\mathrm{if}~b\neq 0,(f^{*}_{1}(b),f^{*}_{2}(b),f^{*}_{3}(b))\in\{(\mathrm{i},\mathrm{i},\mathrm{i}),\\ &(1+\mathrm{i},\mathrm{i},1+\mathrm{i}),(\mathrm{i},1+\mathrm{i},1+\mathrm{i})\}\\ (-1)^{\mathrm{i}}3\cdot 2^{t-1},&\mathrm{if}~b\neq 0,(f^{*}_{1}(b),f^{*}_{2}(b),f^{*}_{3}(b))=(\mathrm{i},\mathrm{i},1+\mathrm{i}),\end{array}\right$.
where $\mathrm{i}\in\{0,1\}$. Clearly, the only case violating the condition $(\mathbf{c}_{2})$ is that for all $b\in{\mathbb{F}}_{2^{m}}^{*}$, $(f^{*}_{1}(b),f^{*}_{2}(b),f^{*}_{3}(b))=(0,0,1)$, which cannot hold in this proposition due to the fact that $f_{1}^{*}$ is bent and then there exists some elements $b$ such that $f^{*}_{1}(b)=1$. Thus $f$ satisfies the conditions $(\mathbf{c}_{1})$ and $(\mathbf{c}_{2})$ in Definition 5.1. ∎
@@
In the rest of this section, we always assume that $m=2t$, ${\rm Tr}(\cdot)={\rm Tr}_{2^{t}}(\cdot)$ for convenience.
We now provide explicit constructions based on Proposition 5.3 and a special subclass of the MM class.
@@
Let $h_{\lambda}(x,y)={\rm Tr}(\lambda x\pi(y))$, where $\lambda\in{\mathbb{F}}_{2^{t}}^{*}$ and $\pi$ is a linear permutation on ${\mathbb{F}}_{2^{t}}$. Let $\lambda_{1},\lambda_{2}$ be distinct elements in $\mathbb{F}_{2^{t}}^{*}$, and $w_{1},w_{2}\in\mathbb{F}^{*}_{2^{t}}$ with $w_{1}\neq w_{2}$. Define some (vectorial) Boolean functions as follows.
$f(x,y)=$ $~h_{\lambda_{1}}(x,y)h_{\lambda_{2}}(x,y),$ (25)
$g(x,y)=$ $~h_{\lambda_{3}}(x,y)+((x+w_{1})^{2^{t}-1}+(x+w_{2})^{2^{t}-1})(y^{2^{t}-1}+1),$ (26)
$F(x,y)=$ $~(f(x,y),g(x,y)),$ (27)
where $\lambda_{3}=\lambda_{1}+\lambda_{2}$.
After plugging the vectorial Boolean function $F$ defined as in (27) into the linear code ${\mathcal{C}}_{F}$ derived in (5), we can find that the resulting linear codes are SO and doubly-even. To obtain singly-even linear codes, we make a slight modification as follows.
${\mathcal{C}}_{F}=\left\{((a_{1},a_{2})\cdot F(x,y)+b_{1}x+b_{2}y)_{(x,y)\in{\mathbb{F}}_{2}^{t}\times{\mathbb{F}}_{2}^{t}\backslash\{(\mathbf{0},\mathbf{0})\}}:(a_{1},a_{2})\in{\mathbb{F}}_{2}\times{\mathbb{F}}_{2},(b_{1},b_{2})\in{\mathbb{F}}_{2^{t}}\times{\mathcal{V}}\right\},$ (28)
where ${\mathcal{V}}=\{x\in{\mathbb{F}}_{2^{t}}~|~{\rm Tr}(xw_{1})={\rm Tr}(xw_{2})=0\}$. It is trivial that $\mathcal{V}$} is a subspace of ${\mathbb{F}}_{2^{t}}$ with the dimension t2t-2. That is to say, our code ${\mathcal{C}}_{F}$ is a subspace of that derived in (5). Obviously, when the $b$ in Lemmas 3.6 and 5.2 belongs to a subspace of $\mathbb{F}_{2^{m}}$, the conclusions of these two lemmas still hold.
To compute the weight distribution of ${\mathcal{C}}_{F}$ derived in (28), it suffices to compute the partial Walsh transform of $W_{F}$, where $b_{2}$ belongs to $\mathcal{V}$}, not ${\mathbb{F}}_{2^{t}}$. We first compute the partial Walsh transforms of the Boolean functions $f$ and $g$ defined as in Definition 5.5.
@@
Let Boolean functions $f$ and $g$ be defined as in Definition 5.5. Let ${\mathcal{V}}=\{x\in{\mathbb{F}}_{2^{t}}~|~{\rm Tr}(xw_{1})={\rm Tr}(xw_{2})=0\}$. Then for any $(b_{1},b_{2})\in{\mathbb{F}}_{2^{t}}\times{\mathcal{V}}$, we have
  1. (1)
    W_f(b_1,b_2) = {2m-1+2t-1, if  (b1,b2)=(0,0),2t-1((-1)A1+ (-1)A2-(-1)A3), if  (b1,b2)(0,0);
  2. (2)
    W_g(b_1,b_2) = (-1)^A_32^t-4;
  3. (3)
    W_f+g(b_1,b_2) = {-2m-1+32t-1-4, if  (b1,b2)=(0,0),2t-1((-1)A1+ (-1)A2+(-1)A3)-4, if  (b1,b2)(0,0),
where for all $i\in\{1,2,3\}$, ${\mathcal{A}}_{i}={\rm Tr}(b_{2}\pi^{-1}(\lambda_{i}^{-1}b_{1}))$.
Proof.
(1) By Lemma 2.3, for any $(b_{1},b_{2})\in{\mathbb{F}}_{2^{t}}\times{\mathcal{V}}$, we have
$W_{f}(b_{1},b_{2})=\left\{\begin{array}[]{ll}2^{m-1}+2^{t-1}\left((-1)^{h_{\lambda_{1}}^{*}(0,0)}+(-1)^{h_{\lambda_{2}}^{*}(0,0)}-(-1)^{h_{\lambda_{3}}^{*}(0,0)}\right),&\text{if}~(b_{1},b_{2})=(0,0),\\ 2^{t-1}\left((-1)^{h_{\lambda_{1}}^{*}(b_{1},b_{2})}+(-1)^{h_{\lambda_{2}}^{*}(b_{1},b_{2})}-(-1)^{h_{\lambda_{3}}^{*}(b_{1},b_{2})}\right),&\text{if}~(b_{1},b_{2})\neq(0,0).\end{array}\right$.
Moreover, since $h_{\lambda_{1}}^{*}(0,0)=h_{\lambda_{2}}^{*}(0,0)=h_{\lambda_{3}}^{*}(0,0)=0$ and $h_{\lambda_{i}}^{*}(b_{1},b_{2})={\mathcal{A}}_{i}$ for all $i\in\{1,2,3\}$ due to Eq. (3), the desiring result is obtained.
(2) By the definition of the Walsh transform, for any $(b_{1},b_{2})\in{\mathbb{F}}_{2^{t}}\times{\mathcal{V}}$, we have
[…]
$=$ $\sum_{(x,y)\in{\mathbb{F}}_{2^{t}}^{2}}(-1)^{h_{\lambda_{3}}(x,y)+{\rm Tr}(b_{1}x+b_{2}y)}-2\left((-1)^{{\rm Tr}(b_{2}w_{1})}+(-1)^{{\rm Tr}(b_{2}w_{2})}\right)$
$=$ $(-1)^{{\mathcal{A}}_{3}}2^{t}-4,$
where the lase equality holds due to Eqs. (1), (3) and $b_{2}\in{\mathcal{V}}$.
(3) Similar to the proof of the above statement, for any $(b_{1},b_{2})\in{\mathbb{F}}_{2^{t}}\times{\mathcal{V}}$, we have
$W_{f+g}(b_{1},b_{2})$ $=$ $\sum_{(x,y)\in{\mathbb{F}}_{2^{t}}^{2}}(-1)^{h_{\lambda_{1}}(x,y)h_{\lambda_{2}}(x,y)+h_{\lambda_{3}}(x,y)+{\rm Tr}(b_{1}x+b_{2}y)}-4$
$=$ $\sum_{(x,y)\in{\mathbb{F}}_{2^{t}}^{2}}(-1)^{h_{\lambda_{1}}(x,y)h_{\lambda_{2}}(x,y)+h_{\lambda_{1}}(x,y)+h_{\lambda_{2}}(x,y)+{\rm Tr}(b_{1}x+b_{2}y)}-4,$
where the second equality holds due to $\lambda_{3}=\lambda_{1}+\lambda_{2}$. Moreover, by Lemma Corollary 2.4, the desiring result can be obtained. ∎
In the following, we first prove that the Boolean functions $f$ and $g$ in Definition 5.5 satisfy the conditions in Definition 5.1.
@@
Let $f$ and $g$ be defined as in Definition 5.5. Then $f$ and $g$ satisfy the conditions in Definition 5.1, where $G=g$, $m=2t,k=2$.
@@
On one hand, note that $f=h_{\lambda_{1}}h_{\lambda_{2}}$, where $h_{\lambda_{1}}={\rm Tr}(\lambda_{1}x\pi(y))$ and $h_{\lambda_{2}}={\rm Tr}(\lambda_{2}x\pi(y))$ belongs to the MM class. By Eq. (3), we have that $h_{\lambda_{1}}={\rm Tr}(\lambda_{1}^{-1}y\pi^{-1}(x))$ and $h_{\lambda_{2}}^{*}={\rm Tr}(\lambda_{2}^{-1}y\pi^{-1}(x))$. It follows from Proposition 5.3 (2) that the Boolean function $f$ defined in Eq. (25) satisfies the conditions $(\mathbf{c}_{1})$ and $(\mathbf{c}_{2})$ in Definition 5.1.
@@
$g(x,y)=\begin{cases}{\rm Tr}(\lambda_{3}x\pi(y)),&~\text{if $(x,y)\notin\{(0,w_{1}),(0,w_{2})\}$}\\ {\rm Tr}(\lambda_{3}x\pi(y))+1,&~\text{if $(x,y)\in\{(0,w_{1}),(0,w_{2})\}$.}\end{cases}$
Thus $g(0,0)=0$ and it is clear that $vf+ug$ is not affine for any nonzero $(v,u)\in{\mathbb{F}}_{2}\times{\mathbb{F}}_{2}$. From Lemma 5.6 (2) and Lemma 2.11, it follows that $\mathcal{C}_{g}$ is a minimal code. ∎
We now compute the partial Walsh transforms of the vectorial Boolean function $F(x,y)$ defined as in Definition 5.5.